# Cascading and Proxy Membership Inference Attacks

### v1

@@ II Problem Definition and Threat Models
    The challenger selects two sets: a subset $D_{\mathrm{a}}\subseteq D$ and a set $D_{\mathrm{b}}$ sampled from $\mathbb{D}$. These two sets are combined to create a query set: $D_{\mathrm{query}}=D_{\mathrm{a}}\cup D_{\mathrm{b}}$, which the challenger then sends to the adversary.
@@
Missed Opportunities of Existing MIAs. Our paper is motivated by the observations that existing MIAs did not fully take advantage of the available information in both adaptive and non-adaptive settings.
$\bullet$ Membership Dependencies in the Adaptive Setting. xisting adaptive MIAs [9, 7, 10] typically predict the membership of each query instance independently. They thus fail to take advantage of the conditional membership dependencies among instances, leading to suboptimal performance.
@@ III Cascading Membership Inference Attack
[…]
@@ III-A Theoretical Intuition of Joint MIA
[…]
@@
Joint MIA with Gibbs Sampling (accurate but prohibitively expensive). In order to sample from the joint distribution $\Pr(\mathbf{M}|o_{\theta})$, it requires capturing the dependencies between the membership indicators introduced by conditioning on $o_{\theta}$. The traditional Gibbs sampling procedure updates each variable sequentially [20], with each update based on the conditional distribution of the variable given the current states of all other variables. This iterative process can be expressed formally at any iteration $t\geq 0$ as:
@@
We now provide a theoretical result to show that it is possible to perform this joint sampling and converge to the true target stationary distribution $\Pr(\mathbf{M}|o_{\theta})$, as well as the almost sure convergence of any metric of success of the attack.
Theorem 1 (Convergence of Joint MIA Gibbs Sampling).
  1. 1.
  2. 2.
    $S_{T}=\frac{1}{T}\sum_{t=1}^{T}L(\mathbf{M}^{(t)},D)$
The proof of Theorem 1 is provided in Section A-A and is based on the stationarity of Markov chains and the convergence properties of martingale theory.
Theorem 1 provides a rigorous foundation for the proposed Gibbs sampling approach, confirming its suitability for Joint MIA by establishing both the convergence to the target distribution and the consistency of performance metrics. However, using this approach is computationally prohibitive, as convergence usually requires many iterations. The challenge now lies in devising a practical and effective method for performing a joint membership inference attack (Joint MIA), which benefits from jointly inferring all instances in $D_{\text{query}}$ without paying the high cost of full Gibbs sampling.
@@ III-B $\mathsf{CMIA}$: A Fast Approximation of Joint MIA
The sampling process proceeds iteratively, with each step refining the ordering of the remaining instances to prioritize those with the highest membership probabilities. Specifically, for the $i$-th sample, the instances are reordered such that $\Pr(M_{i}=1|M_{-i}^{(1,0)},o_{\theta})\geq\Pr(M_{j}=1|M_{-j}^{(1,0)},o_{\theta})$ for $i<j$, where $i=2,\ldots,n$ denotes the current iteration. Ties are again resolved arbitrarily. At each iteration, the process checks if the membership probability for the next instance exceeds a predetermined threshold, in which case the instance is deemed a member, and its membership status is set to $M_{i}=1$. Otherwise, the process terminates and the remaining instances are deemed non-members.
This greedy approach identifies a subset of highly probable members while avoiding unnecessary computations for less likely candidates. Empirically, as we see in Section V-B, joint MIA achieves significantly higher attack performance compared to independent inference of prior work.
@@ III-C Implementation Details of $\mathsf{CMIA}$
4:  ${}\mathcal{P}^{k}_{\text{shadow}}\leftarrow\{\}$triangleright initialize shadow dataset-model set
5:  for $N$ times do
6:   $D_{\text{tmp}}\sim D_{\text{adv}}^{\text{adapt}}$triangleright sample a dataset
@@
11:  ${}\mathcal{S}^{k}\leftarrow\{\}$triangleright initialize membership scores set
@@
Implementation Details. We set the number of the cascading iterations $K=10$ for all experiments, as we observe that all evaluated attacks exhibit significant improvements within only a few iterations (as shown in Section V-B). The stopping criterion $\delta$ is set to 303030, meaning that the iteration halts if fewer than 303030 new anchors can be selected. Although the threshold selection procedure can be repeated by using different ground-truthed shadow models to obtain an average threshold estimate, we find that it remains robust even with a single randomly chosen shadow model. All above hyperparameters are kept fixed across all evaluated attacks and datasets to demonstrate the robustness of the framework.
@@ IV-A Theoretical Intuition of Marginal MIA
[…]
@@ IV-B Attack Method
3:for $N$ times do
4:  $D_{\text{shadow}}\sim D_{\text{adv}}^{\text{non-adapt}}$triangleright sample a shadow dataset
@@ V-A Experimental Setup
  • LOSS [15] uses the loss of the query instance as the score.
V-B Evaluation of $\mathsf{CMIA}$
@@
Main Results. We apply $\mathsf{CMIA}$ with six SOTA MIAs across four image datasets. As shown in Table I, $\mathsf{CMIA}$ consistently boosts all attacks across datasets, particularly in the low false-positive rate regime. For example, $\mathsf{CMIA}$ enhances the performance of LiRA by 5×5$\times$5 × at a false-positive rate of 0.001% and boosts the Calibration attack by 7×7$\times$7 × at the same FPR on MNIST. These results represent a substantial advancement, as such datasets were previously considered difficult to attack. In addition, we find that $\mathsf{CMIA}$ can elevate previously weak attacks to a level comparable to strong ones. For instance, $\mathsf{CMIA}$ boosts the Calibration attack from 0.19% to 0.55% at a false-positive rate of 0.1% on MNIST, nearly matching the performance of LiRA (i.e., 0.56% at the same FPR). The Receiver Operating Characteristic (ROC) curve [40] and results on other model architectures are reported in Section C-A.
(a) MNIST
(b) CIFAR-100
V-C Evaluation of $\mathsf{PMIA}$
@@ V-D Ablation Study
TABLE IV: Impact of the hyperparamters $K$ and $\delta$ in $\mathsf{CMIA}$.
@@
Inference Cost/seconds 1.231.231.23 2.522.522.52 1.851.851.85 3.033.033.03 10.4710.4710.47 >400,000>400,\!000> 400 , 000 49.549.549.5 31.531.531.5 $\mathbf{15.8}$
@@ V-E Additional Investigations
Attack on Non-image Datasets. We also evaluate the attack performance on two non-image datasets to demonstrate the generality of the proposed attacks. Specifically, we train multilayer perceptrons (MLPs) on the Purchase and Texas datasets [1]. For $\mathsf{CMIA}$, we use LiRA as the base model and apply the cascading framework. For $\mathsf{PMIA}$, we employ the Wasserstein distance to select the top 10 proxy models for each query instance. The attack performance under both adaptive and non-adaptive settings is presented in Table VII. As shown, both $\mathsf{CMIA}$ and $\mathsf{PMIA}$ consistently outperform existing attacks, particularly in TPR at low FPR.
@@ Appendix A Missing Proofs
A-A Proof of Theorem 1
@@
We prove this theorem by establishing that the Gibbs sampling procedure forms a Markov chain that satisfies the conditions for the Martingale Convergence Theorem [64].
@@ B-B Accuracy of the Trained Target Model
Model MNIST FMNIST C-10 C-100
Train Val Train Val Train Val Train Val
ResNet50 100.0% 99.5% 100.0% 91.3% 99.9% 90.1% 99.9% 67.9%
VGG16 100.0% 99.5% 100.0% 92.9% 99.9% 87.1% 99.9% 60.8%
DenseNet121 100.0% 99.5% 99.9% 93.3% 99.9% 89.3% 99.9% 63.4%
MobileNetV2 100.0% 99.4% 100.0% 92.4% 99.9% 88.0% 99.9% 61.0%
@@
Model MNIST FMNIST C-10 C-100
Train Val Train Val Train Val Train Val
ResNet50 100.0% 99.5% 100.0% 94.8% 99.7% 90.4% 92.3% 66.9%
VGG16 100.0% 99.7% 100.0% 95.6% 99.9% 91.7% 99.6% 72.3%
DenseNet121 100.0% 99.6% 100.0% 96.1% 99.9% 90.2% 99.9% 71.5%
MobileNetV2 100.0% 99.5% 100.0% 94.9% 99.9% 97.1% 100.0% 72.4%
@@ C-A Experiments of $\mathsf{CMIA}$ on Other Models
TABLE XX: Performance comparison of non-adaptive attacks on DenseNet121 trained on four images datasets.

### v2

@@ II Problem Definition and Threat Models
    The challenger randomly selects a subset $D_{\mathrm{a}}\subseteq D$ and samples a set $D_{\mathrm{b}}$ from $\mathbb{D}$. These two sets are combined to create a query set: $D_{\mathrm{query}}=D_{\mathrm{a}}\cup D_{\mathrm{b}}$, which the challenger then sends to the adversary.
@@
Missed Opportunities of Existing MIAs. Our paper is motivated by the observations that existing MIAs do not fully take advantage of the available information in both adaptive and non-adaptive settings.
$\bullet$ Membership Dependencies in the Adaptive Setting. Existing adaptive MIAs [9, 7, 10] typically predict the membership of each query instance independently. They thus fail to take advantage of the conditional membership dependencies among instances, leading to suboptimal performance.
@@ III Cascading Membership Inference Attack
Figure 1: Statistical dependencies of joint MIA shows that conditioning on output $o_{\theta}$ creates a collider dependency between the membership indicators $M_{i}$ and $M_{j}$, i.e., $M_{i}\not\!\!\bot\!\!\!\bot M_{j}\mid o_{\theta}$.
@@ III-A Theoretical Intuition of Joint MIA
We assume that instances in the size-$n$ query set $D_{\text{query}}$ are ordered in a canonical order, i.e., $D_{\text{query}}=\left\{(x_{i},y_{i})\mid 1\leq i\leq n\right\}$. Given target model $f_{\theta}$ trained using dataset $D$, we define the (unobserved) vector of membership variables as $\mathbf{M}=(M_{1},M_{2},\ldots,M_{n})$, where each $M_{i}$ is a Bernoulli variable, $M_{i}:=\mathbbm{1}[(x_{i},y_{i})\in D]$. We use $o_{\theta}=\{(x_{i},f_{\theta}(x_{i}))|\forall(x_{i},y_{i})\in D_{\text{query}}\}$ to denote the output softmax probabilities that the adversary can obtain from $D_{\text{query}}$.
Existing MIAs [9, 7, 14] typically treat each query instance $(x_{i},y_{i})$ in isolation, computing individual membership probabilities $\Pr(M_{i}=1|o_{\theta})$ independently. At first glance, this approach appears reasonable, as the data instances in the training set $D$ could be assumed to be independently and identically distributed (i.i.d.), implying that the membership indicators $M_{i}$ and $M_{j}$ are marginally independent, i.e., $M_{i}\bot\!\!\!\bot M_{j}$, for all $i\neq j$. However, this independence assumption is violated when conditioning on the model’s output $o_{\theta}$ over the entire query dataset $D_{\text{query}}$. In this conditional setting, the membership indicators, although independent marginally, become dependent due to the collider effect induced by $o_{\theta}$, as illustrated in Figure 1. Specifically, conditioning on $o_{\theta}$ renders $M_{i}$ and $M_{j}$ conditionally dependent, i.e., $M_{i}\not\!\!\bot\!\!\!\bot M_{j}\mid o_{\theta}$.
@@
Joint MIA with Gibbs Sampling (accurate but prohibitively expensive). Sampling from the joint distribution $\Pr(\mathbf{M}|o_{\theta})$ requires capturing the dependencies between the membership indicators introduced by conditioning on $o_{\theta}$. The traditional Gibbs sampling procedure updates each variable sequentially [20], with each update based on the conditional distribution of the variable given the current states of all other variables. This iterative process can be expressed formally at any iteration $t\geq 0$ as:
@@
In Section A-A, we provide a theorem to show that it is possible to perform this joint sampling and converge to the true target stationary distribution $\Pr(\mathbf{M}|o_{\theta})$, as well as the almost sure convergence of any metric of success of the attack.
However, using Gibbs sampling for Joint MIA is computationally prohibitive, as convergence usually requires many iterations. The challenge now lies in devising a practical and effective method for performing a joint membership inference attack, which benefits from jointly inferring all instances in $D_{\text{query}}$ without paying the high cost of full Gibbs sampling.
@@ III-B $\mathsf{CMIA}$: A Fast Approximation of Joint MIA
The sampling process proceeds iteratively, with each step refining the ordering of the remaining instances to prioritize those with the highest membership probabilities. Specifically, for the $i$-th sample, the instances are reordered such that $\Pr(M_{i}=1|M_{-i}^{(1,0)},o_{\theta})\geq\Pr(M_{j}=1|M_{-j}^{(1,0)},o_{\theta})$ for $i<j$, where $i=2,\ldots,n$ denotes the current iteration. Ties are again resolved arbitrarily. At each iteration, the process checks if the membership probability for the next instance exceeds a predetermined threshold, in which case the instance is deemed a member, and its membership status is set to $M_{i}=1$. Otherwise, the process terminates, and the remaining instances are deemed non-members.
This greedy approach identifies a subset of highly probable members while avoiding unnecessary computations for less likely candidates. As such, it can be viewed as a computationally efficient heuristic for Maximum a Posteriori (MAP) estimation, which aims to find the single most likely point from the distribution. Empirically, as we see in Section V-B, joint MIA achieves significantly higher attack performance compared to independent inference of prior work.
@@ III-C Implementation Details of $\mathsf{CMIA}$
Implementation Details. We set the number of the cascading iterations $K=10$ for all experiments, as we observe that all evaluated attacks exhibit significant improvements within only a few iterations (as shown in Section V-B). The stopping criterion $\delta$ is set to $30$, meaning that the iteration halts if fewer than $30$ new anchors can be selected. Although the threshold selection procedure can be repeated by using different ground-truthed shadow models to obtain an average threshold estimate, we find that it remains robust even with a single randomly chosen shadow model. All above hyperparameters are kept fixed across all evaluated attacks and datasets to demonstrate the robustness of the framework.
@@
4:  $\mathcal{P}^{k}_{\text{shadow}}\leftarrow\{\}$ $\triangleright$ initialize shadow dataset-model set
5:  for NN times do
6:   $D_{\text{tmp}}\sim D_{\text{adv}}^{\text{adapt}}$ $\triangleright$ sample a dataset
@@
11:  $\mathcal{S}^{k}\leftarrow\{\}$ $\triangleright$ initialize membership scores set
@@ IV-A Theoretical Intuition of Marginal MIA
A detailed proof based on Bayes’ rule is provided in Section A-B. Note that the expectations in the formula are taken with respect to datasets sampled uniformly at random from $\mathcal{S}^{+}_{(x_{i},y_{i})}$ and $\mathcal{S}^{-}_{(x_{i},y_{i})}$, which are the countable collections of all possible datasets that include or exclude the query instance $(x,y)$. The theorem establishes that the membership posterior odds test $\mathcal{A}_{\mathrm{odds}}$ for an adversary involves comparing a special likelihood ratio to a threshold derived from prior probabilities.
@@
3:for NN times do
4:  $D_{\text{shadow}}\sim D_{\text{adv}}^{\text{non-adapt}}$ $\triangleright$ sample a shadow dataset
@@
IV-B Attack Method
$\phi(f(x)_{y})=\log\left(\frac{f(x)_{y}}{1-f(x)_{y}}\right),$
where $f(x)_{y}$ denotes the softmax probability (aka. confidence score) of the model $f$ on the instance $(x,y)$. This scaling transformation stabilizes the distribution and, empirically, allows it to be well-approximated by a normal distribution. We then adopt the same likelihood estimators used in LiRA [9]:
Overview. The workflow of $\mathsf{PMIA}$ is outlined in Algorithm 2, consisting of two phases: the prepare and inference phases. In the prepare phase, we follow the standard shadow training technique [9, 1] and train $N$ shadow models by randomly sampling shadow datasets from $D_{\text{adv}}^{\text{non-adapt}}$ (lines 3–7). In the inference phase, we first collect the distribution of confidence scores for the target instance $(x,y)$ when it is not in the model’s training set (lines 11–13). Next, we select proxy data from $D_{\text{adv}}^{\text{non-adapt}}$ for the query instance and gather their confidence score distributions when they are in the shadow models’ training sets (lines 15–22). Finally, we estimate the mean and variance for both collected confidence distributions (lines 23–24), query the target model $f_{\theta}$ on $(x,y)$ to output a parametric likelihood ratio $\tilde{\Lambda}$ as the membership score.
@@ V-A Experimental Setup
  • LOSS [15] uses the loss of the query instance as the score.
@@
V-B Evaluation of $\mathsf{CMIA}$
Main Results. We apply $\mathsf{CMIA}$ with six SOTA MIAs across four image datasets. As shown in Table I, $\mathsf{CMIA}$ consistently boosts all attacks across datasets, particularly in the low false-positive rate regime. For example, $\mathsf{CMIA}$ enhances the performance of LiRA by $5\times$ at a false-positive rate of 0.001% and boosts the Calibration attack by $7\times$ at the same FPR on MNIST. These results represent a substantial advancement, as such datasets were previously considered difficult to attack. In addition, we find that $\mathsf{CMIA}$ can elevate previously weak attacks to a level comparable to strong ones. For instance, $\mathsf{CMIA}$ boosts the Calibration attack from 0.19% to 0.55% at a false-positive rate of 0.1% on MNIST, nearly matching the performance of LiRA (i.e., 0.56% at the same FPR). The Receiver Operating Characteristic (ROC) curve [40] and results on other model architectures are reported in Section C-A.
(a) MNIST
(b) CIFAR-100
V-C Evaluation of $\mathsf{PMIA}$
@@ V-D Ablation Study
TABLE IV: Impact of the hyperparameters $K$ and $\delta$ in $\mathsf{CMIA}$.
@@
Inference Cost/seconds $1.23$ $2.52$ $1.85$ $3.03$ $10.47$ >400,000>400,\!000 $49.5$ $31.5$ $\mathbf{15.8}$
@@ V-E Additional Investigations
Attack on Non-image Datasets. We also evaluate the attack performance on two non-image datasets to demonstrate the generality of the proposed attacks. Specifically, we train multilayer perceptrons (MLPs) on the Purchase and Texas datasets [1]. For $\mathsf{CMIA}$, we use LiRA as the base model and apply the cascading framework. For $\mathsf{PMIA}$, we employ the Wasserstein distance to select the top 10 proxy models for each query instance. The attack performance under both adaptive and non-adaptive settings is presented in Table VII. As shown, both $\mathsf{CMIA}$ and $\mathsf{PMIA}$ consistently outperform existing attacks, particularly in TPR at low FPR. It is worth mentioning that the performance of all evaluated MIAs is notably lower on these non-image datasets compared to image datasets such as CIFAR-100, despite both having 100 classes. We attribute this to the smaller generalization gap in models trained on non-image datasets. For instance, as shown in Table XV, the Purchase model achieves training and validation accuracies of 0.96 and 0.80, respectively. In contrast, the CIFAR-100 model shows a much larger gap, with a training accuracy of 0.92 and a validation accuracy of only 0.66. Since MIAs typically exploit overfitting for attack [9], they are less effective against better-generalized models, as observed in the non-image datasets.
@@ Appendix A Missing Proofs
A-A Convergence of Joint MIA Gibbs Sampling
We now provide a theoretical result to show that it is possible to perform the joint sampling and converge to the true target stationary distribution $\Pr(\mathbf{M}|o_{\theta})$, as well as the almost sure convergence of any metric of success of the attack.
Theorem 2 (Convergence of Joint MIA Gibbs Sampling).
  1. 1.
  2. 2.
    $S_{T}=\frac{1}{T}\sum_{t=1}^{T}L(\mathbf{M}^{(t)},D)$
@@ A-A Convergence of Joint MIA Gibbs Sampling
We prove the above theorem by establishing that the Gibbs sampling procedure forms a Markov chain that satisfies the conditions for the Martingale Convergence Theorem [64].
@@ B-B Accuracy of the Trained Target Model
Model MNIST FMNIST C-10 C-100 Purchase Texas
Train Val Train Val Train Val Train Val Train Val Train Val
ResNet50 1.00 0.99 1.00 0.91 0.99 0.90 0.99 0.67 - - - -
VGG16 1.00 0.99 1.00 0.92 0.99 0.87 0.99 0.60 - - - -
DenseNet121 1.00 0.99 0.99 0.93 0.99 0.89 0.99 0.63 - - - -
MobileNetV2 1.00 0.99 1.00 0.92 0.99 0.88 0.99 0.61 - - - -
MLP - - - - - - - - 0.97 0.79 0.96 0.78
@@
Model MNIST FMNIST C-10 C-100 Purchase Texas
Train Val Train Val Train Val Train Val Train Val Train Val
ResNet50 1.00 0.99 1.00 0.94 0.99 0.90 0.92 0.66 - - - -
VGG16 1.00 0.99 1.00 0.95 0.99 0.91 0.99 0.71 - - - -
DenseNet121 1.00 0.99 1.00 0.96 0.99 0.90 0.99 0.72 - - - -
MobileNetV2 1.00 0.99 1.00 0.94 0.99 0.97 1.00 0.72 - - - -
MLP - - - - - - - - 0.96 0.80 0.96 0.79
@@ C-A Experiments of $\mathsf{CMIA}$ on Other Models
TABLE $X$: Performance comparison of non-adaptive attacks on DenseNet121 trained on four images datasets.